OpenStudy (anonymous):

g(x)=x^2+4x+1 How do you turn that in a parabola equation?

OpenStudy (anonymous):

it's already a parabola equation. A "parabola" is a second order polynomial

OpenStudy (anonymous):

Then how do I put it in the form y-k=a(x-h)^2

OpenStudy (anonymous):

Ok, first step is to find out the roots of this equation

OpenStudy (anonymous):

set x^2+4x+1=0

OpenStudy (anonymous):

do you know how to find the roots?